# Pseudo-Asynchronous Local SGD: Robust and Efficient Data-Parallel Training

### v1

@@ B.1 Settings
All experimental results, unless otherwise noted, refer to the hyper parameter configuration with the best results for that metric.The results of the ablation study are shown in Section D. The target loss shown in the training curve plots is based on the loss achieved by DDP within the predefined epoch budget.
@@
    Both the 8M and 125M models were trained following the training protocol described below. We trained the model for 15 epochs with a local batch size of 512 per GPU, and use Cluster A in the experiments with 4 GPUs, and Cluster B in the experiments with 8 GPUs. The global batch size was set to 2048. The inner optimizer learning rate was fixed at 0.001, and we employed AdamW (Loshchilov, 2017) for the inner optimizer with gradient clipping enabled. Regarding outer optimization for DiLoCo and PALSGD, we use Nesterov Momentum SGD with the outer learning rate fixed at 0.1 to 0.2. The synchronization interval $H$ was set to 16, the probabilistic synchronization parameter p = 0.1. The variants of the Local SGD algorithm started after 1024 iterations. For ablation study, the synchronization interval $H$ is set to 32 to 256, the probabilistic synchronization parameter p = 0.025 to 0.5 and $\eta_{t}$ is 0.25 to 64.
  • CIFAR10 Training on VGG-16:
Appendix C Additional Results
@@ C.1 ImageNet-1K on ResNet50
Figure 5 presents the training accuracy curves for ResNet-50 on ImageNet-1K with $K=4$ workers and a synchronization interval of $H=64$, conducted on Cluster C. Consistent with our validation accuracy results, PALSGD is 15.4% faster than DDP and 4.3% faster than DiLoCo, while all methods converge to the same final accuracy of 74.0%.
Figure 5: Training Accuracy of ImageNet-1K on ResNet-50 Experiments ($K=4$ / $H=64$): PALSGD demonstrates faster training compared to DDP and DiLoCo.
@@ C.2.1 GPT-Neo-8M Experiment with 4 GPUs (K=4)
We conducted an additional experiment using GPT-Neo-8M with 4 workers on cluster A. Figure 6 presents both the training and validation loss curves for this setup. Consistent with previous findings, PALSGD achieved the fastest convergence, reaching the target loss significantly earlier than both DiLoCo and DDP. Specifically, PALSGD reduced training time by 23.0% compared to DDP, while DiLoCo provided a more modest 7.9% improvement over DDP. DDP exhibited the slowest convergence, taking the longest time to reach the target loss. While DiLoCo was faster than DDP, it failed to reach the target loss within the given training time.
Figure 6: GPT-Neo Experiments (K=4 / H=16): Training time comparison across distributed algorithm to achieve target training loss (top) and validation loss (bottom). PALSGD achieves fastest convergence and lowest loss, while DDP is slowest and DiLoCo did not achieve target loss.
@@ C.2.2 GPT-Neo-8M Experiment with 8 GPUs (K=8)
Figure 7 presents the training loss curve for the GPT-Neo experiment with 8 workers on Cluster B. Consistent with our validation loss results, PALSGD demonstrates the fastest convergence, reaching the target training loss 20.8% faster than DDP. Among the three methods, DDP converges the slowest, while DiLoCo does not achieve the target loss within the given training time.
@@
Figure 7: GPT-Neo Experiments (K=8 / H=64) on DGX-1 (8 V100 GPUs Connected by NVLINK): Training time comparison across distributed algorithm to achieve target loss.
@@ C.3 TinyStories on GPT-Neo-125M
Figure 8: GPT-Neo-125M Experiments (K=8 / H=16) Training time comparison across distributed algorithm to achieve target training loss.
Figure 8 compares the training time required by PALSGD, DiLoCo, and DDP to reach a common target training. PALSGD and DDP eventually achieve the target loss, whereas DiLoCo fails to reach this target. Furthermore, PALSGD converges significantly faster than DDP. Specifically, PALSGD achieves the target training loss 26.4% faster than DDP. These results highlight PALSGD’s superior efficiency in accelerating training while maintaining model quality.
@@ C.4 CIFAR-10 on VGG-16
We conducted experiments on the CIFAR-10 dataset using the VGG16 architecture Simonyan & Zisserman (2014) on cluster A with $k=4$ workers. Figure 9 presents the training loss and validation accuracy of the DDP, DiLoCo, and PALSGD algorithms for varying synchronization intervals ($H$). As expected, the results show an inverse relationship between $H$ and regularization strength for DiLoCo and PALSGD—i.e., training loss tends to be lower when using DDP or smaller values of $H$. The relatively poor performance of DDP is likely due to its use of large batch sizes. In contrast, DiLoCo and PALSGD operate with batch sizes reduced by a factor of four, which may result in a stronger implicit regularization effect. Compared to DiLoCo, PALSGD consistently achieves lower training loss and higher validation accuracy, suggesting better generalization performance. Moreover, PALSGD is less sensitive to the choice of $H$, showing only a $0.058\%$ drop in accuracy when increasing $H$ from 32 to 1024. These results demonstrate the effectiveness of PALSGD in reducing communication overhead without compromising model performance.
@@
Figure 9: CIFAR-10 with VGG16 Experiments: (Top) Comparison of training loss with respect to H (Sync Interval); (Bottom) Comparison of validation accuracy with respect to H. PALSGD achieves lowest loss and highest accuracy consistently.
Appendix D Ablation Studies
@@
Here, we present the results of an ablation study on hyperparameters using GPT-Neo-8M to train on the TinyStories dataset with K=8. We conducted ablation studies on the following hyperparameters: seeds, optimizer selection, $\eta$ , $p$, and $H$. The seeds represent the error bars in our experiments. While optimizer selection and $\eta$ influence the final loss and accuracy, they have minimal impact on training time, so we focused on evaluating these parameters using accuracy and loss metrics. On the other hand, $p$ and $H$ affect training time, so we report the time taken in addition to accuracy and loss. The result of the ablation study for $H$ is already included in Section 7.3.1
@@ D.1 Sensitivity of Seeds
Table 1: TinyStories on GPT-Neo-8M/ Mean and Variance of Final Validation Loss for PALSGD with Different Optimizer Combination
D.2 Optimizer Selection
@@
Figure 10: Comparison of inner and outer optimizer with GPT-Neo Experiments on Cluster B (K=8, H=64). ‘sgd_m’ stands for SGD with Momentum
D.3 $\eta$ Selection
@@
Figure 11: Comparison of $\eta$ on GPT-Neo Experiments on Cluster B (K=8, H=64)
D.4 $p$ Selection
@@
Figure 12: Comparison of $p$ / GPT-Neo Experiments on Cluster B (K=8, H=64)

### v2

@@ 9 Fully Connected input: 1024, output: 10
All experimental results, unless otherwise noted, refer to the hyper parameter configuration with the best results for that metric. The results of the ablation study are shown in Section E. The target loss shown in the training curve plots is based on the loss achieved by DDP within the predefined epoch budget.
  • Small CNN on CIFAR-10 (Preliminary Simulation Experiments): The number of workers ranged from 4 to 64, while synchronization intervals tested were 32. PALSGD was set with $p$ at 0.25. The outer optimizer used was Nesterov momentum, and the inner optimizer was momentum SGD. Inner learning rates of 0.01, 0.001, and 0.0001 were explored. Additionally, $\eta$ was fixed at 1, and outer learning rate included 1, 0.1, and 0.01. No weight decay was applied, and training was conducted for 100 epochs.
  • CIFAR10 Training on VGG-16:
@@
    Both the 8M and 125M models were trained following the training protocol described below. We trained the model for 15 epochs with a local batch size of 512 per GPU, and use Cluster A in the experiments with 4 GPUs, and Cluster B in the experiments with 8 GPUs. The global batch size was set to 2048. The inner optimizer learning rate was fixed at 0.001, and we employed AdamW (Loshchilov, 2017) for the inner optimizer with gradient clipping enabled. Regarding outer optimization for DiLoCo and PALSGD, we use Nesterov Momentum SGD with the outer learning rate fixed at 0.1 to 0.2. The synchronization interval $H$ was set to 16 (125M) or 64 (8M), the probabilistic synchronization parameter $p=0.1$, $\eta_{t}$ = 16. The variants of the Local SGD algorithm started after 1024 iterations. For ablation study, the synchronization interval $H$ is set to 32 to 256, the probabilistic synchronization parameter $p$ = 0.025 to 0.5 and $\eta_{t}$ is 0.25 to 64.
Appendix C Full Results of Experiments
@@ C.1 ImageNet-1K on ResNet50
Figure C.1 presents the training accuracy curves for ResNet-50 on ImageNet-1K with $K=4$ workers and a synchronization interval of $H=64$, conducted on Cluster C. Consistent with our validation accuracy results, PALSGD is 15.4% faster than DDP and 4.3% faster than DiLoCo, while all methods converge to the same final accuracy of 74.0%.
Figure C.1: Training Accuracy of ImageNet-1K on ResNet-50 Experiments ($K=4$ / $H=64$): PALSGD demonstrates faster training compared to DDP and DiLoCo.
@@ C.2.1 GPT-Neo-8M Experiment with 4 GPUs (K=4)
We conducted an additional experiment using GPT-Neo-8M with 4 workers on cluster A. Figure C.2 presents both the training and validation loss curves for this setup. Consistent with previous findings, PALSGD achieved the fastest convergence, reaching the target loss significantly earlier than both DiLoCo and DDP. Specifically, PALSGD reduced training time by 23.0% compared to DDP, while DiLoCo provided a more modest 7.9% improvement over DDP. DDP exhibited the slowest convergence, taking the longest time to reach the target loss. While DiLoCo was faster than DDP, it failed to reach the target loss within the given training time.
Figure C.2: GPT-Neo Experiments (K=4 / H=16): Training time comparison across distributed algorithm to achieve target training loss (top) and validation loss (bottom). PALSGD achieves fastest convergence and lowest loss, while DDP is slowest and DiLoCo did not achieve target loss.
@@ C.2.2 GPT-Neo-8M Experiment with 8 GPUs (K=8)
Figure C.3 presents the training loss curve for the GPT-Neo experiment with 8 workers on Cluster B. Consistent with our validation loss results, PALSGD demonstrates the fastest convergence, reaching the target training loss 20.8% faster than DDP. Among the three methods, DDP converges the slowest, while DiLoCo does not achieve the target loss within the given training time.
@@
Figure C.3: GPT-Neo Experiments (K=8 / H=64) on DGX-1 (8 V100 GPUs Connected by NVLINK): Training time comparison across distributed algorithm to achieve target loss.
@@ C.3 TinyStories on GPT-Neo-125M
Figure C.4: GPT-Neo-125M Experiments (K=8 / H=16) Training time comparison across distributed algorithm to achieve target training loss.
Figure C.4 compares the training time required by PALSGD, DiLoCo, and DDP to reach a common target training. PALSGD and DDP eventually achieve the target loss, whereas DiLoCo fails to reach this target. Furthermore, PALSGD converges significantly faster than DDP. Specifically, PALSGD achieves the target training loss 26.4% faster than DDP. These results highlight PALSGD’s superior efficiency in accelerating training while maintaining model quality.
@@ C.4 CIFAR-10 on VGG-16
We conducted experiments on the CIFAR-10 dataset using the VGG16 architecture Simonyan & Zisserman (2014) on cluster A with $k=4$ workers. Figure C.5 presents the training loss and validation accuracy of the DDP, DiLoCo, and PALSGD algorithms for varying synchronization intervals ($H$). As expected, the results show an inverse relationship between $H$ and regularization strength for DiLoCo and PALSGD—i.e., training loss tends to be lower when using DDP or smaller values of $H$. The relatively poor performance of DDP is likely due to its use of large batch sizes. In contrast, DiLoCo and PALSGD operate with batch sizes reduced by a factor of four, which may result in a stronger implicit regularization effect. Compared to DiLoCo, PALSGD consistently achieves lower training loss and higher validation accuracy, suggesting better generalization performance. Moreover, PALSGD is less sensitive to the choice of $H$, showing only a $0.058\%$ drop in accuracy when increasing $H$ from 32 to 1024. These results demonstrate the effectiveness of PALSGD in reducing communication overhead without compromising model performance.
@@
Figure C.5: CIFAR-10 with VGG16 Experiments: (Top) Comparison of training loss with respect to H (Sync Interval); (Bottom) Comparison of validation accuracy with respect to H. PALSGD achieves lowest loss and highest accuracy consistently.
Appendix D Convergence Analysis by Global Steps
D.1 Image Classification Tasks: ImageNet-1K on ResNet 50
Figure D.1: Training (Top) and Validation (Bottom) Accuracy of ImageNet-1K on ResNet-50 Experiments ($K=4$ / $H=64$) with using Cluster C.
Our main results in Figure 2 present wall-clock time comparisons. To isolate the source of performance gains, this section provides a complementary analysis. We compare methods based on the number of global steps required to reach a target metric. This approach distinguishes algorithmic convergence speed from system-level speedups due to reduced communication. Image Classification on ImageNet-1K For the ResNet-50 experiment, all three methods—PALSGD, DiLoCo, and DDP—reached a given validation accuracy in a nearly identical number of global steps, as shown in Figure D.1 (bottom). DDP demonstrated a marginal advantage in the convergence of training accuracy. These results suggest the convergence properties of the algorithms are comparable for this task. The practical wall-clock advantage of PALSGD shown in Figure 2 therefore stems primarily from its reduction in communication overhead, not from superior step-wise convergence.
D.2 Language Modeling Tasks: TinyStories on GPT-Neo
D.2.1 GPT-Neo-8M Experiments
The GPT-Neo-8M experiments show that PALSGD and DDP again achieve similar convergence for both training and validation loss. In contrast, DiLoCo failed to converge, with its loss increasing during training. PALSGD’s advantage over DDP on this model is its communication efficiency. Its advantage over DiLoCo is its stable convergence.
Figure D.2: GPT-Neo 8M Experiments ($K=8$ / $H=64$) on DGX-1 (8 V100 GPUs Connected by NVLINK): Training time comparison across distributed algorithm to achieve target loss.
D.2.2 GPT-Neo-125M Experiments
Experiments with the larger GPT-Neo-125M model reveal a different dynamic. Here, PALSGD achieves the target loss in fewer steps than both DDP and DiLoCo. PALSGD required 15.8% fewer steps than DDP to reach the target validation loss. DiLoCo again failed to converge. This algorithmic improvement contributes significantly to the overall performance gain. The 24.4% wall-clock time reduction reported in Figure 3 is a product of both faster algorithmic convergence and reduced communication costs.
Figure D.3: GPT-Neo 125M Experiments ($K=8$ / $H=16$) on DGX-1 (8 V100 GPUs Connected by NVLINK): Training time comparison across distributed algorithm to achieve target loss.
Appendix E Ablation Studies
Here, we present the results of an ablation study on hyperparameters using GPT-Neo-8M to train on the TinyStories dataset with K=8. We conducted ablation studies on the following hyperparameters: seeds, optimizer selection, $\eta$ , $p$, and $H$. Default hyperparameter setting is described in Appendix B.
@@ Appendix E Ablation Studies
The seeds represent the error bars in our experiments. While optimizer selection and $\eta$ influence the final loss and accuracy, they have minimal impact on training time, so we focused on evaluating these parameters using accuracy and loss metrics. On the other hand, $p$ and $H$ affect training time, so we report the time taken in addition to accuracy and loss. The result of the ablation study for $H$ is already included in Section 7.4.
@@ E.1 Sensitivity of Seeds
Table E.1: TinyStories on GPT-Neo-8M/ Mean and Variance of Final Validation Loss for PALSGD with Different Optimizer Combination
E.2 Optimizer Selection for Loss
@@
Figure E.1: Comparison of inner and outer optimizer with GPT-Neo Experiments on Cluster B (K=8, H=64). ‘sgd_m’ stands for SGD with Momentum
E.3 Optimizer Selection for Convergence
Figure E.2 compares the time required for different outer optimizers to reach the target validation loss. The results show that the Nesterov optimizer achieves the shortest training time. In contrast, using SGD as the outer optimizer is significantly slower than using either Nesterov or SGD Momentum. Notably, when the outer optimizer is SGD, our algorithm reduces to a randomized variant of Local SGD.
Figure E.2: Comparison of inner and outer optimizer with GPT-Neo Experiments on Cluster B (K=8, H=64).
E.4 $\eta$ Selection
@@
Figure E.3: Comparison of $\eta$ on GPT-Neo Experiments on Cluster B (K=8, H=64)
E.5 $p$ Selection
@@
Figure E.4: Comparison of $p$ / GPT-Neo Experiments on Cluster B (K=8, H=64)
Appendix F Additional Study
F.1 CIFAR10 Experiments on Scaling Performance
In this section, we conduct distributed training experiments using multiple nodes on Cluster C, rather than the simulated environment described in the previous section. PALSGD, DiLoCo, and DDP were evaluated using the CIFAR-10 dataset on the VGG-16 architecture. Experiments were executed on Cluster C with configurations of 1, 2, 4 and 8 nodes corresponding to 4, 8, 16 and 32 GPUs, respectively.
For optimization, local models utilized SGD with momentum. The outer optimizer was Nesterov momentum. Workers synchronized parameters every 128 steps. The local SGD phase began after 2048 global iterations, except for the K=16, 32 setting, where it started after 1024, 512 global iterations.
A grid search determined key hyperparameters. Inner learning rates were selected from {0.025, 0.05, 0.075}. Outer learning rates ranged across {0.2, 0.4, 0.8, 1.2, 1.6}. Algorithm-specific parameters included a gradient-drop probability $p$ selected from 0.02, 0.05, 0.1 and an eta coefficient chosen from {0.1, 0.25, 0.5}.
Several parameters remained constant. The batch size per GPU was set at 256, with no learning rate decay applied throughout training. To ensure high computational efficiency, we used the largest possible per-GPU batch size. The total number of training epochs was fixed for all experiments. This setup means that as the number of GPUs increases, the global batch size grows proportionally, while the total number of training steps decreases.
Figure F.1: Comparison of $K$ / CIFAR10 Experiments on Cluster C (K=4, 8, 16, 32, H=128)
Our baseline, DDP, showed a 4% drop in validation accuracy when scaled from 4 to 16 GPUs. The degradation was more severe at 32 GPUs, where accuracy fell by over 15%. This sharp decline on 32 GPUs is likely attributable to the large-batch problem on CIFAR-10. The global batch size becomes 8,192, which means an entire epoch over the 50,000-image dataset completes in just over six steps. Training time for all methods decreased linearly on a log scale, indicating good system scalability. PALSGD and DiLoCo have nearly identical runtimes. Their advantage in training time becomes more significant when scaling from 16 to 32 GPUs. This result suggests that both methods effectively mitigate the communication bottleneck that emerges at larger scales. However, they also suffer from accuracy degradation at 32 GPUs. Despite this, both methods consistently achieve higher validation accuracy than DDP. More importantly, PALSGD demonstrates superior accuracy to DiLoCo across all scales.